A student sets up the foll ation to solve a proble Full steps and answer. Thanks.
Answer #1

0.57 mol(1ml/10^-3L) / (2.3mol/L) = 0.57 mol(1ml/1ml) / (2.3mol/L)

                                                 = 0.57mol*L/ 2.3mol = 0.57L/2.3 = 0.247 L